OpenStudy (inowalst):

What is the simplified form of \[\frac{ 1 }{ x }-\frac{ 2 }{ x^2-x }\]

OpenStudy (inowalst):

I think its \[\frac{ x-3 }{ x(x-1) }\] Because 1-2 is -3. Then x times x is x^2. Then x times -1 is -x.

OpenStudy (inowalst):

I'm not completely sure.

OpenStudy (aum):

correct.

OpenStudy (inowalst):

:o Really?

OpenStudy (aum):

\[ \frac{1}{x} - \frac{2}{x^2-x} = \frac{1}{x} - \frac{2}{x(x-1)} = \frac{1}{x}*\frac{x-1}{x-1} - \frac{2}{x(x-1)} = \\ \frac{x-1}{x(x-1)} - \frac{2}{x(x-1)} = \frac{x-1-2}{x(x-1)} = \frac{x-3}{x(x-1)} \]

OpenStudy (inowalst):

Wow thank you!

OpenStudy (aum):

you are welcome.

OpenStudy (theraggedydoctor):

dot.